OpenStudy (anonymous):

A child is creating a pyramid with building blocks. The top three levels include 3 blocks, 7 blocks, and 11 blocks. Part 1: How many blocks would be needed for a pyramid 25 levels tall? 2: Use complete sentences to explain how a sum of an arithmetic series was applied.

OpenStudy (anonymous):

every level has 4 more then the one above, so by using this formula (3+0 X 4) (3+1 X 4) and so on i can see, i dont know where to go from here

OpenStudy (mathmate):

What does n represent in your formula N=(3+4n)? What would N be for n=1,2,...

OpenStudy (anonymous):

@mathmate how come u adjust 25 in this formula

OpenStudy (mathmate):

I am using the formula proposed by OP, and suggested some verifications to make adjustments, if required.

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

does n = 25?

OpenStudy (anonymous):

i guess sum is 25.not sure though.

OpenStudy (mathmate):

The formula you proposed works like this: when n=1, N=3+4(1)=7, which corresponds to the number of blocks for the second level, and when n=2, N=3+4(2)=11, which corresponds to the third level. If you want the formula to accept n as the level number, you will need to adjust it as follows: N=3+4(n-1)=4n-1 Check N with this new formula for n=1,2 and 3 before you attempt the 25th level.

OpenStudy (anonymous):

so the for a 25 level pyramid the bottom level would have 103 blocks? do i have to do this for all 25 levels or is there a formula/shortcut?

OpenStudy (mathmate):

Did you check the formula for n=1,2 and 3? If it worked for 1,2 and 3, chances are good that it works for 25 (directly).

OpenStudy (anonymous):

it works for my original formula, the formula you mentioned is just a more complicated formula with more steps imo

OpenStudy (anonymous):

this is an arithmetic series... an = a1 + (n - 1)d a25 = 3 + (25 - 1) * 4 a25 = 3 + 24 * 4 a25 = 3 + 96 a25 = 99 -- this is how many blocks that are on the 25th level sn = n(a1 + an)/2 s25 = 25(3 + 99) / 2 s25 = 25(102) / 2 s25 = 2550/2 s25 = 1275 -- this is the sum of all blocks all the way to the 25th level